OpenStudy (anonymous):

It takes 8 hours to fill a tank with a particular chemical. Without treatment, all of the chemical in the tank would evaporate in 12 hours. If the tank is empty to start, and the chemical is not being treated as it enters the tank, how long will it take to have a full tank?

OpenStudy (mayankdevnani):

24 hours In this one the tank is filling at 1/3 the rate so will take 3 times as long to fill

OpenStudy (mayankdevnani):

ok @stupidinmath

OpenStudy (mayankdevnani):

@mathslover am i right?

mathslover (mathslover):

http://answers.yahoo.com/question/index?qid=20080108172709AAMgx5j Seems so :)

OpenStudy (anonymous):

this was my guess too. thankyou so much @mayankdevnani :D

OpenStudy (mayankdevnani):

oh!!!!!!! @stupidinmath very-very welcome

OpenStudy (phi):

this is a rate * time = distance (but distance is 1 filled tank) rate to fill is 1 tank per 8 hours or 1/8 rate to empty is 1/12 the difference between filling and emptying is 1/8 - 1/12 = 1/24 now do rate * time= 1 tank 1/24 * t =1 t= 24 hours

OpenStudy (anonymous):

Oh. Nice explanation, @phi . thanks also:)